# Expert-Guided Diffusion Planner for Auto-bidding

### v1

@@ 4.1. Blended Forcing with Optimal Expert
To bridge this gap, we propose a Variational Autoencoder (VAE)-enhanced hybrid mechanism called Blended Forcing, which implicitly models environmental dynamics through the lens of expert historical behaviors. The VAE’s latent space is trained to encode the distribution of optimal expert actions conditioned on environmental states. During inference, the VAE reconstructs a pseudo ground-truth action that approximates expert-level decisions given the current state, thereby serving as a dynamically adaptive substitute for the missing ground truth.
Moreover, we apply for coverage $\mathbf{x}^{e}(\tau)$ to $\hat{\mathbf{x}}^{e}(\tau)$ with a hyperparameter $\delta$ to harmonize the representation of expert, which can be formulated as:
@@
where $\mathbb{I}[\cdot]$ represents the indicator function, and in each $k$-th diffusion step, it will choose $TF(\cdot)$ in the probability of $\delta$ and $DF(\cdot)$ otherwise. As for the structure of $DF(\cdot)$, we implement one layer DNN for both the encoder and decoder with the learning parameters $\phi$ for reconstruction.
@@ 4.2. Expert-guided Conditional Planner
(12) $\mathbb{C}=\underbrace{\hat{\mathbf{x}}^{e}(\tau)}_{IMPLICIT}\oplus\quad\underbrace{[f_{\text{}}(R(\tau)\oplus f_{\text{}}^{\prime}(C(\tau))]}_{EXPLICIT(\mathbb{C}^{g})}.$
@@
To formulate the query for cross-attention, we concatenate the original sequence generated in a single step, denoted as $x_{k}(\tau)$, the total number of denoising steps $k$, and the relevant conditions $\mathbb{C}^{g}$ as query.
(15) $\mathbb{O}=x_{k}(\tau)\oplus k\oplus\mathbb{C}^{g},\quad\mathbb{Q}=\text{LN}\left[\text{MLP}[\mathbb{O}]\right]$
It is noted that $\text{LN}(\cdot)$ and $\text{MLP}(\cdot)$ represent the layer normalization and dense layer, respectively, and cross-attention is defined as the process of extracting effective information from the expert sequence as follows.
(16) $\mathrm{Cross\_Attn}=\mathrm{Softmax}\left(\frac{(\mathbb{Q}W^{Q})(\mathbb{C}W^{K})^{T}}{\sqrt{d}}\right)(\mathbb{C}W^{V}),$
the Key and Value of Cross-Attention is defined in (12) and the parameter matrices $W^{Q}$$W^{K}$$W^{V}$ represent the weights associated with the transformations of the original $Q$,$K$, and $V$, and $d$ denotes the dimension after the vectors have been concatenated. The latent representation $\mathbb{H}_{0}$ can be obtained from cross-attention, and $x_{k-1}(\tau)$ is obtained as follows.
(17) $\mathbb{H}_{1}=\text{FFN}\left[\text{LN}[\mathbb{H}_{0}+\mathbb{O}]\right],\ \text{and}\ x_{k-1}(\tau)=\text{MLP}\left(\mathbb{H}_{1}+\mathbb{H}_{0}\right),$
where FFN is Feed Forward Network, and $\mathbb{H}_{0}$ and $\mathbb{H}_{1}$ are add as output.
@@ 4.3. Efficient Action Generation
4.3.2. Joint Optimization.
As for optimization, considering that VAE is employed for extracting information from expert sequences, the VAE loss (Kingma, 2013) must also be incorporated as $\mathcal{L}_{\text{exp}}$, which primarily consists of the reconstruction loss of $z$ and the constraint that the intermediate latent variable $z$ follows a normal distribution. Moreover, to inverse actions through the series of states generated by the conditional diffusion planner, we utilize the inverse dynamics loss $\mathcal{L}_{\text{inv}}$ to optimize the reconstruction of the action patterns following (Guo et al., 2024b). Therefore, the optimized objective can be simplified as:
@@
where the $\xi$ governs the reconstruction of the expert trajectory and action generation. The overall training and inference phases are shown in Algorithm 1 and 2, respectively.
@@ 5.1. Experimental Setting
We conduct extensive experiments in offline environments to demonstrate the effectiveness of our proposed method. Four research questions are investigated in the following experiments:
@@
We utilize Alibaba [^1] https://github.com/alimama-tech/AuctionNet simulated environment from the NeurIPS 2024 competition (Tianchi, 2024). The dataset covers 21 periods (7-27), each with over 500,000 impression opportunities and 48 steps. Each opportunity involves 48 bidding agents. The dataset contains over 500 million records, each with predicted conversion values, bids, auction details, and impression results. For a fair comparison, data from periods 7-13 is for training and 14-20 for testing.
@@
(20) $\begin{aligned} score=\mathbb{P}(CPA,C)\sum_{i}x_{i}E_{i}V_{i},\quad\mathbb{P}(CPA,C)=\text{min}\{(\frac{CPA}{C})^{\lambda},1\}.\end{aligned}$
$x_{i}$ indicates whether the agent secures the impression, $E_{i}$ signifies if the ad is successfully shown to the user after winning, and $V_{i}$ denotes whether a conversion occurs after the user is exposed to the advertisement. The advertiser sets the target $CPA$, while $C$ represents the actual $CPA$ after completing the process. The penalty function defines a hyperparameter $\lambda$, usually set to 2.
@@
To verify the effectiveness of our proposed method, we list some common reinforcement learning methods as benchmarks for comparison.
  • CQL (Kumar et al., 2020) addresses overestimation of out-of-distribution actions by introducing conservative constraints in the Q-function
  • IQL (Kostrikov et al., 2022) focuses on in-sample actions and avoiding the querying of values for unseen actions
  • BC (Torabi et al., 2018) is trained through supervised learning by directly replicating expert trajectory data
  • DT (Chen et al., 2021) Transformer architecture-based algorithm that maps trajectory data to action outputs through sequence modeling
  • DiffBid (Guo et al., 2024a) conditional diffusion modeling approach for bid generation
@@ 5.2.1. Overall Performances (RQ1)
We compare the overall performance of  EGDB  with various baselines, as reported in Table 1. Our method consistently outperforms other approaches across all dates and achieves the highest scores on each date, validating the effectiveness of our proposed approach. This superiority arises from the robust generative capabilities of the diffusion model and the guidance provided by expert samples, which enhances the model’s ability to generate superior bidding trajectories. The IQL and CQL methods outperform BC, suggesting that traditional reinforcement learning can develop more effective strategies than simple supervised learning. Additionally, DT demonstrates superior performance compared to traditional reinforcement learning methods at certain times, indicating that taking sequence dependency into account is beneficial for optimizing bidding strategies.
@@ 5.2.2. Ablation Studies (RQ2)
To further clarify the contributions of each module in  EGDB , we conduct ablation experiments by evaluating various configurations of  EGDB . Table  2 presents the overall results of our ablation studies on the proposed components, revealing three key insights: (1) All modules contribute positively, with the CA module demonstrating a significant impact; (2) An adaptive fast sampling strategy can achieve optimal performance-accuracy trade-offs; (3) Blended forcing partially emulates expert sequence behaviors, indicating promising regularization potential.
Table 2. The ablation study results for our proposed  EGDB: (1) w/o Acc. $\rightarrow$ the model denoises step by step to replace fast-sampling; (2) w/o BF. $\rightarrow$ blended-forcing is removed and only the auto-encoder; (3) w/o CA. $\rightarrow$ Cross-Attention block is removed, and self-attention is adopted.
@@ 5.2.3. Parameter Sensitivity (RQ3)
We further analyze the sensitivity of hyperparameter settings in  EGDB. As shown in Figure 3, we select the periods with the lowest (P15) and highest (P17) metrics for display to ensure our samples are representative.
  • $\gamma$ : The number of skip steps in accelerated sampling.
  • $\xi$ : balances the weight of the reconstruction of the expert trajectory and action generation loss within the loss function;
  • $\delta$ : A hyperparameter to harmonize the representation of the expert.
The experimental results indicate that the parameters $\gamma$ and $\xi$ are significantly influenced by hyperparameter changes, whereas $\delta$ is relatively less affected. Furthermore, it has been demonstrated that employing an appropriate skipping strategy can yield optimal results. Since a larger value of $\xi$ determines the model’s ability to reconstruct the expert and the action, placing greater emphasis on reconstruction may restrict the generalization capabilities of the Denoising Diffusion Probabilistic Model (DDPM), resulting in a gradual decline in performance metrics.
@@ 5.3. Online A/B Testing(RQ4)
To verify the performance of the proposed method in real-world business scenarios, we selected a batch of the lowest-cost campaigns aimed at maximizing conversions within budget constraints. Diffbid (Guo et al., 2024a) serves as a baseline to evaluate the effectiveness of our proposed method, as our modifications are built upon it. Additionally, the performance of Diffbid on offline datasets is surpassed only by our proposed method. We conducted a budget-splitting and traffic-splitting online experiment from February 7, 2024, to February 13, 2024. The experiment results are presented in Table 3.
  • Feature: a series of window features with different window sizes, including the bid coefficient, remaining traffic, remaining budget, cumulative consumption, cumulative revenue, etc.
  • Action: the difference between two consecutive bid coefficients is represented as $a_{t}=b_{t+1}-b_{t}$, where $b_{t}$ denotes the bid coefficient for a specific advertiser at time $t$
Our method demonstrated a significant improvement compared to the baseline model. Specifically,  EGDB  increased conversions by 11.29% and revenue by 12.35%. In contrast to DiffBid, EGDB  primarily integrates Variational Autoencoder (VAE) expert trajectory generation and cross-attention-based denoising steps, which influence inference complexity. However, we utilized the skip sampling method to reduce 25% of the inference resources when serving the same advertising campaign while significantly maintaining the model’s effectiveness.
Table 3. Performance of  EGDB  in Online Testing.
@@
Baseline (IQL) 269,686 247 265,155 1,188,336
Ours 269,686 247 295,090 1,335,214
compare +11.29% +12.36%
@@ 6. Conclusion
In this paper, we propose EGDB, an expert-guided conditional diffusion model for auto-bidding. To enhance the diffusion planner with a more robust and personalized condition, we extract expert trajectories and introduce the EGCD block, which integrates expert conditions into diffusion transformers. The inherent behavior of the expert is modeled using a Variational Autoencoder (VAE), and we propose a blended forcing technique to improve training efficiency. Additionally, we adopt accelerated sampling with a skip-step approach to speed up the generation process for matching online inference. Offline and online experiments have been designed to demonstrate superiority over baseline models. Future work will explore optimal path generation principles and hybrid architectures that combine diffusion models with the advantages of linear programming.

### v2

@@ 4.1. Blended-Forcing with Optimal Expert
To bridge this gap, we propose a VAE-enhanced hybrid mechanism called Blended-Forcing, which implicitly models environmental dynamics through the lens of expert historical behaviors. The VAE’s latent space is trained to encode the distribution of optimal expert actions conditioned on environmental states. During inference, the VAE reconstructs a pseudo ground-truth action that approximates expert-level decisions given the current state, thereby serving as a dynamically adaptive substitute for the missing ground truth.
Moreover, we apply for coverage $\mathbf{x}^{e}(\tau)$ to $\hat{\mathbf{x}}^{e}(\tau)$ with a hyperparameter $\delta$ to harmonize the representation of the expert, which can be formulated as:
@@
where $\mathbb{I}[\cdot]$ represents the indicator function, and in each $k$-th diffusion step, it will choose $TF(\cdot)$ in the probability of $\delta$ and $DF(\cdot)$ otherwise. As for the structure of $DF(\cdot)$, we implement a one-layer DNN for both the encoder and decoder with the learning parameters $\phi$ for reconstruction.
@@ 4.2. Expert-Guided Conditional Planner
(12) $\mathbb{C}=\underbrace{\hat{\mathbf{x}}^{e}(\tau)}_{IMPLICIT}\oplus\quad\underbrace{[f_{\text{}}(R(\tau)\oplus f_{\text{}}’(C(\tau))]}_{EXPLICIT(\mathbb{C}^{g})}.$
@@
To formulate the query for cross-attention, we concatenate the original sequence generated in a single step, denoted as $x_{k}(\tau)$, the total number of denoising steps $k$, and the relevant conditions $\mathbb{C}^{g}$ as the query:
(15) $\mathbb{O}=x_{k}(\tau)\oplus k\oplus\mathbb{C}^{g},\ \text{and}\ \mathbb{Q}=\texttt{L-Norm}[\text{MLP}(\mathbb{O})].$
It is noted that $\texttt{L-Norm}[\cdot]$ and $\text{MLP}(\cdot)$ represent the layer normalization and dense layer, respectively, and cross-attention is defined as the process of extracting effective information from the expert sequence as follows:
(16) $\text{Cross-Attn}=\texttt{Softmax}\left(\frac{(\mathbb{Q}W^{Q})(\mathbb{C}W^{K})^{T}}{\sqrt{d}}\right)(\mathbb{C}W^{V}),$
where the Key and Value of cross-attention is defined in (12) and the parameter matrices $W^{Q}$$W^{K}$$W^{V}$ represent the weights associated with the transformations of the original $Q$,$K$, and $V$, and $d$ denotes the dimension after the vectors have been concatenated. The latent representation $\mathbb{H}_{0}$ can be obtained from cross-attention, and $x_{k-1}(\tau)$ is obtained as follows:
(17) $\mathbb{H}_{1}=\text{FFN}(\texttt{L-Norm}[\mathbb{H}_{0}+\mathbb{O}]),\ \ x_{k-1}(\tau)=\text{MLP}\left(\mathbb{H}_{1}+\mathbb{H}_{0}\right),$
where $\text{FFN}(\cdot)$ is the Feed-Forward Network, $\mathbb{H}_{0}$ and $\mathbb{H}_{1}$ are add as output.
@@ 4.3. Efficient Action Generation
4.3.2. Final Optimization.
As for optimization, considering that VAE is employed for extracting information from expert sequences, the VAE loss (Kingma, 2013) must also be incorporated as $\mathcal{L}_{\text{exp}}$, which primarily consists of the reconstruction loss of $z$ and the constraint that the intermediate latent variable $z$ follows a normal distribution. Moreover, to inverse actions through the series of states generated by the conditional diffusion planner, we utilize the inverse dynamics loss $\mathcal{L}_{\text{inv}}$ to optimize the reconstruction of the action patterns following (Guo et al., 2024b). Therefore, the optimization can be simplified as:
@@
Here, $\xi$ governs the reconstruction of the expert trajectory and action generation. The overall training and inference phases are shown in Algorithm 1 and Algorithm 2, respectively.
@@ 5.1. Experimental Setting
We conduct extensive experiments both in offline environments and online A/B testing to demonstrate the effectiveness of our proposed method. Four research questions are investigated in the following experiments:
@@
We utilize Alibaba [^1] https://github.com/alimama-tech/AuctionNet simulated environment from the NeurIPS 2024 competition (Tianchi, 2024). The dataset covers 21 periods (7-27), each with over 500,000 impression opportunities and 48 steps. Each opportunity involves 48 bidding agents. The dataset contains over 500 million records, each with predicted conversion values, bids, auction details, and impression results. For a fair comparison, data from periods 7-13 is for training and 14-20 for evaluation.
@@
(20) $\begin{aligned} \text{score}=\mathbb{P}(CPA,C)\sum_{i}x_{i}E_{i}V_{i},\ \text{and}\\ \mathbb{P}(CPA,C)=\text{min}\{(\frac{CPA}{C})^{\lambda},1\}.\end{aligned}$
Here, $x_{i}$ indicates whether the agent secures the impression, $E_{i}$ signifies if the ad is successfully shown to the user after winning, and $V_{i}$ denotes whether a conversion occurs after the user is exposed to the advertisement. The advertiser sets the target $CPA$, while $C$ represents the actual $CPA$ after completing the process. The penalty function defines a hyperparameter $\lambda$, usually set to 2.
@@
To verify the effectiveness of our proposed method, we list some common reinforcement learning methods as benchmarks for comparison:
  • CQL (Kumar et al., 2020) addresses overestimation of out-of-distribution actions by introducing conservative constraints in the Q-function.
  • IQL (Kostrikov et al., 2022) focuses on in-sample actions and avoiding the querying of values for unseen actions.
  • BC (Torabi et al., 2018) is trained through supervised learning by directly replicating expert trajectory data.
  • DT (Chen et al., 2021) is a Transformer-based framework that maps trajectories to action outputs through sequence modeling.
  • DiffBid (Guo et al., 2024a) utilizes conditional diffusion modeling approach for bid generation.
  • EBaReT (Li et al., 2025) integrates the Bag Reward Decision Transformer with expert trajectory generation.
@@ 5.2.1. Overall Performances (RQ1)
We compare the overall performance of  EGDB  with various baselines in Table 1. Our method consistently outperforms others across most periods and ranks second in the rest, validating the effectiveness of our proposed approach. This advantage arises from the robust generative capabilities of the diffusion model and the guidance provided by expert samples, which enhances the model’s ability to generate superior bidding trajectories. The IQL and CQL methods outperform BC, suggesting that traditional reinforcement learning can develop more effective strategies than simple supervised learning. Moreover, DT-based methods (e.g., EBaReT) demonstrate superior performance compared to conventional reinforcement learning methods at specific periods, indicating that taking sequence dependency into account is beneficial for optimizing bidding strategies.
@@ 5.2.2. Ablation Studies (RQ2)
To further clarify the contributions of each module in  EGDB , we conduct ablation experiments by evaluating various configurations of  EGDB. Table 2 presents the overall results of our ablation studies on the proposed components, revealing three key insights: (1) All modules contribute positively, with the CA. Block demonstrating a significant impact; (2) An adaptive fast-sampling strategy can achieve optimal performance-accuracy trade-offs; (3) Blended-Forcing partially emulates expert sequence behaviors, indicating promising regularization potential.
Table 2. The ablation study results for our proposed  EGDB: (1) w/o Acc. $\rightarrow$ the model denoises step by step to replace acceleration; (2) w/o BF. $\rightarrow$ Blended-Forcing is removed and only the Auto-Encoder; (3) w/o CA. $\rightarrow$ Cross-Attention block is removed, and Self-Attention is adopted.
@@ 5.2.3. Parameter Sensitivity (RQ3)
We further analyze the sensitivity of hyperparameter settings in  EGDB. As shown in Figure 3, we select the periods with the lowest (P15) and highest (P17) metrics for display to ensure that our samples are representative. We have chosen the following three hyperparameters, whose detailed descriptions are provided below: (i) $\gamma\Rightarrow$ number of skip steps to accelerate diffusion sampling; (ii) $\xi\Rightarrow$ balancing the weight of the reconstruction of the expert trajectory and action generation within the loss functions; (iii) $\delta\Rightarrow$ harmonizing the expert representation within the blend-forcing paradigm. The experimental results indicate that the parameters $\gamma$ and $\xi$ are significantly influenced by hyperparameter changes, whereas $\delta$ is relatively less affected. Furthermore, it has been demonstrated that employing an appropriate skipping strategy can yield optimal results. Since a larger value of $\xi$ determines the model’s ability to reconstruct the expert and the action, placing greater emphasis on reconstruction may restrict the generalization capabilities of the DDPM, resulting in a gradual decline in performance metrics.
@@ 5.3. Online A/B Testing (RQ4)
To verify the performance of the proposed method in real-world business scenarios, we selected a batch of the lowest-cost campaigns, aimed at maximizing conversions within budget constraints. DiffBid (Guo et al., 2024a) serves as a baseline to evaluate the effectiveness of our proposed method, as our modifications are built upon it. We conducted a budget and traffic-splitting online experiment from 7 February 2025 to 13 February 2025, respectively, and the online features and actions can be detailed in the following:
  • Features: a series of window features with different window sizes, including the bid coefficient, remaining traffic, remaining budget, cumulative consumption, cumulative revenue, etc.
  • Actions: the difference between two consecutive bid coefficients is represented as $a_{t}=b_{t+1}-b_{t}$, where $b_{t}$ denotes the bid coefficient for a specific advertiser at time $t$.
The experiment results are presented in Table 3. Our method demonstrated a significant improvement compared to the baseline model. Specifically,  EGDB  increased conversions by 11.29% and revenue by 12.35%. In contrast to DiffBid, EGDB  primarily integrates VAE expert trajectory generation and cross-attention-based denoising steps, which influence inference complexity. However, we utilized the skip sampling method to reduce 25% of the inference resources when serving the same advertising campaign while significantly maintaining the model’s effectiveness.
Table 3. Performance of  EGDB  in Online A/B Testing.
@@
DiffBid 269,686 247 265,155 1,188,336
EGDB 269,686 247 295,090 1,335,214
Improv. +11.29% +12.36%
@@ 6. Conclusion
In this paper, we propose EGDB, an expert-guided conditional diffusion model for auto-bidding. To enhance the diffusion planner with a more robust and personalized condition, we extract expert trajectories and introduce the EGCD block, which integrates expert conditions into diffusion transformers. The inherent behavior of the expert is modeled using the VAE block, and we propose a blended-forcing technique to improve training efficiency. Additionally, we adopt accelerated sampling with a skip-step approach to speed up the generation process for matching online inference. Offline and online experiments have been designed to demonstrate superiority over baseline models. Future work will explore optimal path generation principles and hybrid architectures that combine diffusion models with the advantages of linear programming.